OpenStudy (anonymous):

how do you solve for x: (x+4)/2=7 ?

OpenStudy (anonymous):

What is your understanding of this question? We can go from there :)

OpenStudy (anonymous):

well im not sure how to deal with fractions when solving for x

OpenStudy (anonymous):

Also.. \(\Large\cal\color{magenta}{Welcome~to~OpenStudy!}\)

OpenStudy (anonymous):

thnks lol

OpenStudy (anonymous):

In order to do this, you must apply the inverse operation to get rid of the fraction. Do you know how?

OpenStudy (anonymous):

multiply the fraction by 2?

OpenStudy (anonymous):

Yes! awesome job :)

OpenStudy (anonymous):

ok so now we're left with 2x+8=7 or 2x+8=14?

OpenStudy (anonymous):



OpenStudy (anonymous):

or neither ?

OpenStudy (anonymous):

oh ok i got it now

OpenStudy (anonymous):

Try again :) the 2 will cancel out

OpenStudy (anonymous):

so x+4=14 then substract 4 on both sides were left with x=10?

OpenStudy (anonymous):